IS WATER ICE THE PRECURSOR TO OH^{+} AND H_{2}O^{+} IN ORION KL?

\underline{PAUL B. RIMMER}, Department of Physics, The Ohio State University, Columbus, OH 43210; ERIC HERBST, Departments of Astronomy, Chemistry and Physics, The Ohio State University, Columbus, OH 43210.

The reactive ions OH^{+} and H_{2}O^{+} have been observed in an outflow in front of the Orion KL region at significant column densities of ∼ 10^{13} cm^{−2} with the Heterodyne Instrument for the Far Infrared (HIFI) on the Herschel Space Observatory. No H_{3}O^{+} was observed, establishing an upper limit of ≈ 10^{12} cm^{−2}. This is unexpected, because both OH^{+} and H_{2}O^{+} react with molecular hydrogen to form H_{3}O^{+}. The primary destruction of H_{3}O^{+} is by recombination with electrons. We explore the low velocity Orion KL outflow with a gas-grain PDR model where UV radiation, cosmic rays, X-rays, and temperature depend on both depth into the cloud and time. The model starts with cold core conditions and a radiation field of χ = 1 and ζ_{H2} = 5 × 10^{−17} s^{−1} at the edge. Water ice collects on the grains at this time, and then as stars form, χ increases to 10^{4} and ζ_{H2} becomes 5 × 10^{−15} s^{−1} at the edge. At all times, temperature is calculated via thermal balance using the Meudon PDR code. At A_{V} < 4 into the cloud the water desorbs off grains and becomes ionized by cosmic rays and X-rays, and dissociated by UV photons, increasing the rates of OH^{+} and H_{2}O^{+} formation. On the other hand, the increased electron fraction depletes the H_{3}O^{+}. The results of this model agree to within a factor of 5 with observation, and place the H_{3}O^{+} column at ≈ 5 × 10^{11} cm^{−2}. We will discuss the model and its results for the OH^{+} and H_{2}O^{+} ions as well as predicted abundances for other species.